# Transic: Sim-to-Real Policy Transfer by Learning from Online Correction

### v1

@@ 1 Introduction
Learning in simulation is a potential approach to the realization of generalist robots capable of solving sophisticated decision-making tasks [1, 2]. Learning to solve these tasks requires a large amount of training data [3, 4, 5]. Providing unlimited training supervision [6] through state-of-the-art simulation [7, 8, 9, 10, 11] could alleviate the burden of collecting data in the real world with physical robots [12, 13]. Therefore, it is crucial to seamlessly transfer and deploy robot control policies acquired in simulation, usually through reinforcement learning (RL), to real-world hardware. Successful demonstrations of this simulation-to-reality (sim-to-real) approach have been shown in dexterous in-hand manipulation [14, 15, 16, 17, 18], quadruped locomotion [19, 20, 21, 22], biped locomotion [23, 24, 25, 26, 27, 28], and quadrotor flight [29, 30].
Nevertheless, replicating similar success in manipulation tasks with robotic arms remains surprisingly challenging, with only a few cases in simple non-prehensile manipulation (such as pulling, pushing, and pivoting objects) [31, 32, 33, 34], industry assembly under restricted settings [35, 36, 37, 38, 39], drawer opening [40], and peg swinging [40]. The difficulty mainly stems from the unavoidable sim-to-real gaps [41, 11], including but not limited to perception gap [42, 43, 19, 44], embodiment mismatch [45, 19, 46], controller inaccuracy [47, 48, 49], and dynamics realism [50]. Traditionally, researchers tackle these sim-to-real gaps and the transferring problem through system identification [51, 19, 52, 31], domain randomization [53, 14, 54, 55], real-world adaptation [56, 57], and simulator augmentation [58, 59, 60]. Many of these approaches require explicit, domain-specific knowledge and expertise on tasks or simulators. Although for a particular simulation-reality pair, there may exist specific inductive biases that can be hand-crafted post hoc to close the sim-to-real gap [19], such knowledge is often not available a priori. Identifying its effects on task completion is also intractable.
@@
Our key insight is that the human-in-the-loop framework is promising for addressing the sim-to-real gaps as a whole, in which humans directly assist the physical robots during policy execution by providing online correction signals. The knowledge required to close sim-to-real gaps can be learned from human signals. To this end, we present Transic (transferring policies sim-to-real by learning from online correction, Fig. 1), a data-driven approach to enable successful transferring of robot manipulation policies trained with RL in simulation to the real world. In Transic, once the base robot policies are acquired from simulation training, they are deployed onto real robots where human operators monitor the execution. When the robot makes mistakes or gets stuck, humans interrupt and assist robot policies through teleoperation. Such human intervention data are collected to train a residual policy, after which the base policy and the residual policy are combined to solve contact-rich manipulation tasks, such as furniture assembly. With the synergetic integration with previous approaches, since humans can successfully assist the robot trained in silico to complete real-world tasks, sim-to-real gaps are implicitly handled and addressed by humans in a domain-agnostic manner. Additionally, human supervision naturally guarantees safe deployment.
(a) Simulation Policy
(b) Direct Deployment
(c) Human-in-the-Loop Correction
(d) Successful Transfer with Learned Residual Policy
Figure 1: Transic for sim-to-real transfer in contact-rich robotic manipulation tasks. a) and b) Naïvely deploying policies trained in simulation usually fails due to various sim-to-real gaps. Here, the robot attempts to first align the light bulb with the base and then insert and screw the light bulb into the base. c) A human operator monitors robot behaviors, intervenes, and provides online correction through teleoperation when necessary. Human data are collected to train a residual policy to tackle various sim-to-real gaps in a holistic manner. d) The simulation policy and residual policy are integrated together during test time to achieve successful sim-to-real transfer for contact-rich tasks, such as screwing a light bulb into the base.
@@
To summarize, the key contribution of our work is a novel, holistic human-in-the-loop method called Transic to tackle sim-to-real transfer of policies for manipulation tasks. Through extensive evaluation, we show that our method leads to more effective sim-to-real transfer compared to traditional methods [51, 53] and requires less real-robot data compared to the prevalent imitation learning and offline RL algorithms [66, 67, 68, 69]. We demonstrate that successful sim-to-real transfer of short-horizon skills can solve long-horizon, contact-rich manipulation tasks in our daily activities, such as furniture assembly. Videos and code are available at transic-robot.github.io.
2 Preliminaries
2.1 Problem Formulation
We formulate a robot manipulation task as an infinite-horizon discrete-time Partially Observable Markov Decision Process (POMDP) $\mathcal{M}\coloneqq\left(\mathcal{S},\mathcal{O},\Omega,\mathcal{A},\mathcal{T},R,\gamma,\rho_{0}\right)$, where $\mathcal{S}$ is the state space, $\mathcal{O}$ is the observation space, and $\mathcal{A}$ is the action space. At time step $t$, a robot observes $o_{t}\in\mathcal{O}$ emitted from observation function $\Omega\left(o_{t}|s_{t},a_{t-1}\right):\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{O}$, executes an action $a_{t}$, and receives a scalar reward $r_{t}$ from the reward function $R(s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The environment proceeds to the next state governed by the transition function $\mathcal{T}(s_{t+1}|s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The robot learns a parameterized policy $\pi_{\theta}\left(\cdot|o\right):\mathcal{O}\rightarrow\Delta\mathcal{A}$ to maximize the expected discounted return $\mathcal{J}\coloneqq\mathbb{E}_{\tau\sim p_{\pi_{\theta}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$ over induced trajectory distribution $\tau\coloneqq\left(s_{0},o_{0},a_{0},r_{0},...\right)\sim p_{\pi_{\theta}}$, where $s_{0}\sim\rho_{0}$ is sampled from the initial state distribution. Additionally, $\gamma\in[0,1)$ is a discount factor. In this work, we model simulation and real environments as two different POMDPs.
2.2 Intervention-Based Policy Learning
We adopt an intervention-based learning framework [66, 67, 70] where a human operator can intervene and take control during the execution of the robot base policy $\pi_{B}$. Denote the human policy as $\pi_{H}$, the following combined policy is deployed during data collection: $\pi^{deployed}=\mathds{1}^{H}\pi^{H}+\left(1-\mathds{1}^{H}\right)\pi^{B},$ where $\mathds{1}^{H}$ is a binary function indicating human interventions. Introducing a trajectory distribution $q(\tau)$ that consists of two observation-action distributions generated by the robot $\rho^{B}$ and human operator $\rho^{H}$, the original RL objective leads to the maximization of a variational lower bound on logarithmic return [71, 67]: $\mathcal{J}(\theta,q)=\mathbb{E}_{q(\tau)}\left[\log R(\tau)+\log p_{\pi_{\theta}}-\log q(\tau)\right],$ where $p_{\pi_{\theta}}$ is the induced trajectory distribution. While the human operator optimizes this objective through intervention and correction, the robot learner maximizes it through $\theta=\arg\max_{\theta\in\Theta}\mathbb{E}_{(o,a)\sim q(\tau)}\left[\log\pi_{\theta}(a|o)\right]$. Various intervention-based policy learning methods have been derived by weighting observation-action pairs differently. For example, HG-Dagger [66] completely ignores robot data $\mathcal{D}^{B}$ and only trains on human data $\mathcal{D}^{H}$ that contain intervention samples. This is equivalent to $q(\tau)\propto\rho^{H}$. Intervention Weighted Regression (IWR) [67] balances the data distribution by emphasizing human intervention: $q(\tau)\propto\alpha\rho^{H}+\rho^{B}$ with $\alpha=|\mathcal{D}^{B}|/|\mathcal{D}^{H}|$. Non-intervention-based methods such as traditional behavior cloning (BC) [72] only learn on $\mathcal{D}^{H}$ with full human demonstrations instead of intervention. This effectively sets $q(\tau)\propto\rho^{H}$.
3 Transic: Sim-to-Real Policy Transfer by Learning from Online Correction
@@
Figure 2: Transic method overview. a) Base policies are first trained in simulation through action space distillation with demonstrations generated by RL teacher policies. Base policies take point cloud as input to reduce perception gap. b) After acquiring base policies, they are first deployed where a human operator monitors the execution. The human intervenes and corrects through teleoperation when necessary. Such intervention and correction data are collected to learn residual policies. Finally, both residual policies and base policies are integrated during test time to achieve successful transfer.
An overview of Transic is shown in Fig. 2. At a high level, after training the base policy in simulation, we deploy it on the real robot while monitored by a human operator. The human interrupts the autonomous execution when necessary and provides online correction through teleoperation. Such intervention and online correction are collected to train a residual policy, after which both base and residual policies are deployed to complete contact-rich manipulation tasks. In this section, we first elaborate on the simulation training phase with several important design choices that reduce sim-to-real gaps before transfer. We then introduce residual policies learned from human intervention and online correction. Subsequently, we present an integrated framework for deploying the base policy alongside the learned residual policy during testing. Finally, we provide implementation details.
@@ 3.1 Learning Base Policies in Simulation with RL
Object geometry matters for contact-rich manipulation. For example, a robot should ideally insert a light bulb into the lamp base with the thread facing down. To retain such 3D information and facilitate sim-to-real transfer, we propose to use point cloud as the main visual modality. Typical RGB observation used in visuomotor policy training [73] suffers from several drawbacks that hinder successful transfer, such as the vulnerability to different camera poses [74] and discrepancies between synthetic and real images [43]. Well-calibrated point cloud observation can bypass these issues and has been successfully demonstrated [75, 15]. During the simulation RL training phase, we synthesize point cloud observations for higher throughput. Concretely, given the synthetic point cloud of the $m$-th object $\mathbf{P}^{(m)}\in\mathbb{R}^{K\times 3}$, we transform it into the global frame through $\mathbf{P}_{g}^{(m)}=\mathbf{P}^{(m)}\left(\mathbf{R}^{(m)}\right)^{\intercal}+\left(\mathbf{p}^{(m)}\right)^{\intercal}$. Here $\mathbf{R}^{(m)}\in\mathbb{R}^{3\times 3}$ and $\mathbf{p}^{(m)}\in\mathbb{R}^{3\times 1}$ denote the object’s orientation and translation in the global frame. Further, the point cloud representation of a scene $\mathbf{S}$ with $M$ objects is aggregated as $\mathbf{P}^{\mathbf{S}}=\bigcup_{m=1}^{M}\mathbf{P}_{g}^{(m)}$ and subsequently used as policy input.
@@
A suitable action abstraction is critical for efficient learning [47, 48] as well as sim-to-real transfer [49]. A high-level controller such as the operational space controller (OSC) [76] facilitates RL exploration [47] but may hinder sim-to-real transfer because it requires accurate modeling of robot parameters, such as joint friction, mass, and inertia [77]; on the other hand, a low-level action space such as the joint position ensures consistent deployment in simulation and real hardware, but renders trial-and-error RL impractical. We draw inspiration from the teacher-student framework [78, 79, 80, 16] and propose to first train the teacher policy $\pi^{teacher}$ through RL with OSC and then distill successful trajectories into the student policy $\pi^{student}$ with joint position control. Specifically, we roll out $\pi^{teacher}$ and record the robot’s joint position at every simulated time interval $\delta t$ to construct a dataset $\mathcal{D}^{teacher}=\{\tau^{(n)}\}_{n=1}^{N}$. We then relabel actions from the end-effector’s poses to joint positions. Such a relabeled dataset is ready to train student policies through behavior cloning. We name this approach as action space distillation and find it crucial to overcome the sim-to-real controller gap. Furthermore, teacher policies directly receive privileged observations for ease of learning, while student policies learn on synthetic point-cloud inputs to match real-world measurements. The student policy parameterized by $\theta$, $\pi^{student}_{\theta}$, is trained by minimizing the following loss function:
$\mathcal{L}^{student}=-\mathbb{E}_{\mathcal{D}^{teacher}}\left[\log\pi^{student}_{\theta}\right]+\beta\mathbb{E}_{\mathcal{D}^{pcd}}\left[\|\phi(\mathbf{P}^{real})-\phi(\mathbf{P}^{sim})\|^{2}\right],$ (1)
where $\phi(\cdot)$ denotes the point cloud encoder of $\pi_{\theta}^{student}$ and $\mathcal{D}^{pcd}=\{(\mathbf{P}^{real},\mathbf{P}^{sim})^{(i)}\}_{i=1}^{N}$ is a separate dataset that contains $N$ pairs of matched point clouds in simulation and reality for regularization purpose. We empirically justify the importance of such regularization in Sec. 4.5.
@@ 3.2 Learning Residual Policies from Online Correction
Once the student policy is obtained from simulation, it is directly used as the base policy $\pi^{B}$ to bootstrap the data collection. Naïvely deploying the base policy on real robots usually results in inferior performance and unsafe motion due to various sim-to-real gaps. In Transic, the base policy is instead deployed in a way that is fully synchronized with the human operator. Concretely, at time step $t$, once $a_{t}^{B}\sim\pi^{B}$ is deployed, the execution is paused and a human operator needs to decide whether intervention is necessary, indicated as $\mathds{1}^{H}_{t}$. Intervention is not necessary for most task execution when the robot is approaching objects of interest. However, when the robot tends to behave abnormally due to unaddressed sim-to-real gaps, the human operator intervenes and takes full control through teleoperation to correct robot errors. In these cases, the robot’s pre- and post-intervention states, as well as intervention indicator $\mathds{1}^{H}_{t}$, are collected to construct the online correction dataset $\mathcal{D}^{H}\leftarrow\mathcal{D}^{H}\cup\left(\mathds{1}^{H}_{t},\mathbf{q}_{t}^{pre},\mathbf{q}_{t}^{post}\right)$. This procedure is illustrated in Algorithm 1.
@@ 3.3 An Integrated Deployment Framework
While in the original formulation of the residual policy, residual actions are always applied [82]. In practice, we find that learning a gating function to predict whether to apply the residual policy or not leads to smoother and more human-like behaviors. We call this learned gated residual policy. Denote the gating function as $g_{\psi}(\cdot)$. It can be trained through classification on $\mathcal{D}^{H}$ and shares the same feature encoder with $\pi^{R}$. The policy effectively being deployed to autonomously complete tasks is an integration of base policy $\pi^{B}$ and residual policy $\pi^{R}$, gated by $g$: $\pi^{deployed}=\pi^{B}\oplus g\pi^{R}$. A joint position controller is used during deployment.

### v2

@@ Abstract
Refer to caption (a) Simulation Policy Refer to caption (b) Direct Deployment Refer to caption (c) Human-in-the-Loop Correction Refer to caption (d) Successful Transfer with Learned Residual Policy
Figure 1: Transic for sim-to-real transfer in contact-rich robotic manipulation tasks. a) and b) Naïvely deploying policies trained in simulation usually fails due to various sim-to-real gaps. Here, the robot attempts to first align the light bulb with the base and then insert and screw the light bulb into the base. c) A human operator monitors robot behaviors, intervenes, and provides online correction through teleoperation when necessary. Human data are collected to train a residual policy to tackle various sim-to-real gaps in a holistic manner. d) The simulation and the residual policies are integrated together during test time to achieve a successful sim-to-real transfer for contact-rich tasks, such as screwing a light bulb into the base.
@@ 1 Introduction
Learning in simulation is a potential approach to the realization of generalist robots capable of solving sophisticated decision-making tasks [1, 2]. Learning to solve these tasks requires a large amount of training data [3, 4, 5]. Providing unlimited training supervision through state-of-the-art simulation [6, 7, 8, 9, 10] could alleviate the burden of collecting data in the real world with physical robots [11, 12]. Therefore, it is crucial to seamlessly transfer and deploy robot control policies acquired in simulation, usually through reinforcement learning (RL), to real-world hardware. Successful demonstrations of this simulation-to-reality (sim-to-real) approach have been shown in dexterous in-hand manipulation [13, 14, 15, 16, 17], locomotion [18, 19, 20, 21, 22, 23, 24, 25, 26, 27], and quadrotor flight [28, 29].
Nevertheless, replicating similar success in manipulation tasks with robotic arms remains surprisingly challenging, with only a few cases in simple non-prehensile manipulation [30, 31, 32, 33], industry assembly under restricted settings [34, 35, 36, 37, 38], and peg swinging [39]. The difficulty mainly stems from the unavoidable sim-to-real gaps [40], including but not limited to perception gap [41, 42, 43], embodiment mismatch [44, 18, 45], controller inaccuracy [46, 47, 48], and dynamics realism [49]. Traditionally, researchers tackle them through system identification [50, 18, 51, 30], domain randomization [52, 13, 53, 54, 55], real-world adaptation [56, 57], and simulator augmentation [58, 59, 60]. Many of these approaches require explicit, domain-specific knowledge and expertise in tasks or simulators. Although for a particular simulation-reality pair, there may exist specific inductive biases that can be hand-crafted post hoc to close the sim-to-real gap [18], this knowledge is often not available a priori. Identifying its effects on task completion is also intractable.
@@
Our key insight is that the human-in-the-loop framework is promising for addressing the sim-to-real gaps as a whole, in which humans directly assist the physical robots during policy execution by providing online correction signals. The knowledge required to close sim-to-real gaps can be learned from human signals. We present Transic (transferring policies sim-to-real by learning from online correction, Fig. 1), a data-driven approach to enable the successful transfer of robot manipulation policies trained with RL in simulation to the real world. In Transic, once the base robot policies are acquired from simulation training, they are deployed to real robots where human operators monitor the execution. When the robot makes mistakes or gets stuck, humans interrupt and assist robot policies through teleoperation. Such human intervention data are collected to train a residual policy, after which the base policy and the residual policy are combined to solve contact-rich manipulation tasks. Unlike previous approaches that heavily rely on domain knowledge, since humans can successfully assist the robot trained in silico to complete real-world tasks, sim-to-real gaps are implicitly handled and addressed by humans in a domain-agnostic manner.
@@
To summarize, our key contribution is a novel, holistic human-in-the-loop method called Transic to tackle sim-to-real policy transfer for manipulation tasks. Through extensive evaluation, we show that our method leads to more effective sim-to-real transfer compared to traditional methods [50, 52] and requires less real-robot data compared to the prevalent imitation learning and offline RL algorithms [66, 67, 68, 69]. We demonstrate that successful sim-to-real transfer of short-horizon skills can solve long-horizon, contact-rich manipulation tasks, such as furniture assembly.
2 Sim-to-Real Policy Transfer by Learning from Online Correction
@@
Figure 2: Transic method overview. a) Base policies are first trained in simulation through action space distillation with demonstrations generated by RL teacher policies. Base policies take point cloud as input to reduce perception gap. b) The acquired base policies are first deployed with a human operator monitoring the execution. The human intervenes and corrects through teleoperation when necessary. Such correction data are collected to learn residual policies. Finally, both residual policies and base policies are integrated during test time to achieve a successful transfer.
An overview of Transic is shown in Fig. 2. This section starts with a brief preliminary review, followed by a description of simulation training. We then introduce residual policies learned from human intervention and online correction and present an integrated framework for deployment during testing. Lastly, we provide implementation details.
2.1 Preliminaries
We formulate a robot manipulation task as an infinite-horizon discrete-time Markov Decision Process (MDP) $\mathcal{M}\coloneqq\left(\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma,\rho_{0}\right)$, where $\mathcal{S}$ is the state space, and $\mathcal{A}$ is the action space. At time step $t$, a robot observes $s_{t}\in\mathcal{S}$, executes an action $a_{t}$, and receives a scalar reward $r_{t}$ from the reward function $R(s_{t},a_{t})$. The environment progresses to the next state following the transition function $\mathcal{T}(s_{t+1}|s_{t},a_{t})$. The robot learns a parameterized policy $\pi_{\theta}\left(\cdot|s\right)$ to maximize the expected discounted return $\mathcal{J}\coloneqq\mathbb{E}_{\tau\sim p_{\pi_{\theta}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$ over induced trajectory distribution $\tau\coloneqq\left(s_{0},a_{0},r_{0},...\right)\sim p_{\pi_{\theta}}$, where $s_{0}\sim\rho_{0}$ is sampled from the initial state distribution and $\gamma\in[0,1)$ is a discount factor. We consider simulation and real environments as two different MDPs. We adopt an intervention-based learning framework [66, 67] where a human operator can intervene and take control during the execution of the robot policy.
@@ 2.2 Learning Base Policies in Simulation with RL
Object geometry matters for contact-rich manipulation. For example, a robot should ideally insert a light bulb into the lamp base with the thread facing down. To retain such 3D information and facilitate sim-to-real transfer, we propose to use point cloud as the main visual modality. Typical RGB observation used in visuomotor policy training [70] suffers from several drawbacks that hinder successful transfer, such as vulnerability to different camera poses [71] and discrepancies between synthetic and real images [42]. Well-calibrated point cloud observation can bypass these issues and has been successfully demonstrated [72, 14]. During the simulation RL training phase, we synthesize point cloud observations for higher throughput. Concretely, given the synthetic point cloud of the $m$-th object $\mathbf{P}^{(m)}\in\mathbb{R}^{K\times 3}$, we transform it into the global frame through $\mathbf{P}_{g}^{(m)}=\mathbf{P}^{(m)}\left(\mathbf{R}^{(m)}\right)^{\intercal}+\left(\mathbf{p}^{(m)}\right)^{\intercal}$. Here, $\mathbf{R}^{(m)}\in\mathbb{R}^{3\times 3}$ and $\mathbf{p}^{(m)}\in\mathbb{R}^{3\times 1}$ denote the object’s orientation and translation in the global frame. Further, the point cloud representation of a scene $\mathbf{S}$ with $M$ objects is aggregated as $\mathbf{P}^{\mathbf{S}}=\bigcup_{m=1}^{M}\mathbf{P}_{g}^{(m)}$ and subsequently used as policy input.
@@
A suitable action abstraction is critical for efficient learning [46, 47] as well as sim-to-real transfer [48]. A high-level controller such as the operational space controller (OSC) [73] facilitates RL exploration [46] but may hinder sim-to-real transfer because it requires accurate modeling of robot parameters, such as joint friction, mass, and inertia [74]; on the other hand, a low-level action space such as the joint position ensures consistent deployment in simulation and real hardware, but renders trial-and-error RL impractical. We draw inspiration from the teacher-student framework [75, 76, 77, 15] and propose to first train the teacher policy $\pi^{teacher}$ through RL with OSC and then distill successful trajectories into the student policy $\pi^{student}$ with joint position control. Specifically, we roll out $\pi^{teacher}$ and record the robot’s joint position at every simulated time interval $\delta t$ to construct a dataset $\mathcal{D}^{teacher}=\{\tau^{(n)}\}_{n=1}^{N}$. We then relabel actions from the end-effector’s poses to joint positions. Such a relabeled dataset is ready to train student policies through behavior cloning. We name this approach as action space distillation and find it crucial to overcome the sim-to-real controller gap. Furthermore, teacher policies directly receive privileged observations for ease of learning, while student policies learn on synthetic point-cloud inputs to match real-world measurements. The student policy parameterized by $\theta$, $\pi^{student}_{\theta}$, is trained by minimizing the loss function: $\mathcal{L}^{student}=-\mathbb{E}_{\mathcal{D}^{teacher}}\left[\log\pi^{student}_{\theta}\right]+\beta\mathbb{E}_{\mathcal{D}^{pcd}}\left[\|\phi(\mathbf{P}^{real})-\phi(\mathbf{P}^{sim})\|^{2}\right]$, where $\phi(\cdot)$ denotes the point cloud encoder of $\pi_{\theta}^{student}$ and $\mathcal{D}^{pcd}=\{(\mathbf{P}^{real},\mathbf{P}^{sim})^{(i)}\}_{i=1}^{N}$ is a separate dataset that contains $N$ pairs of matched point clouds in simulation and reality for regularization purpose. Further justifications of this distillation phase can be found in Appendix D.1.
@@ 2.3 Learning Residual Policies from Online Correction
Once the student policy is obtained from simulation, it is used directly as the base policy $\pi^{B}$ to bootstrap the data collection. Naïvely deploying the base policy on real robots usually results in inferior performance and unsafe motion due to various sim-to-real gaps. In Transic, the base policy is instead deployed in a way that is fully synchronized with the human operator. Concretely, at time step $t$, once $a_{t}^{B}\sim\pi^{B}$ is deployed, a human operator needs to decide whether intervention is necessary, indicated as $\mathds{1}^{H}_{t}$. Intervention is not necessary for most task execution when the robot is approaching objects of interest. However, when the robot tends to behave abnormally, the human operator intervenes and takes full control through teleoperation to correct robot errors. In these cases, the robot’s pre- and post-intervention states, as well as intervention indicator $\mathds{1}^{H}_{t}$, are collected to construct the online correction dataset $\mathcal{D}^{H}\leftarrow\mathcal{D}^{H}\cup\left(\mathds{1}^{H}_{t},\mathbf{q}_{t}^{pre},\mathbf{q}_{t}^{post}\right)$. This procedure is illustrated in Appendix Algorithm 1.
@@ 2.4 An Integrated Deployment Framework
In practice, we find that learning a gating function to control whether to apply residual actions or not leads to better empirical performance (Appendix D.3). We call this learned gated residual policy. Denote the gating function as $g_{\psi}(\cdot)$. It shares the same feature encoder with the residual policy $\pi^{R}$ and is jointly learned through classification on the same correction dataset $\mathcal{D}^{H}$. At the inference time, we effectively use $g_{\psi}$ as an indicator function $\mathbbm{1}_{g}$ to determine whether to apply the predicted residual actions. The policy effectively being deployed to autonomously complete tasks is an integration of base policy $\pi^{B}$ and residual policy $\pi^{R}$, gated by $g$: $\pi^{deployed}=\pi^{B}\oplus\mathbbm{1}_{g}\pi^{R}$. A joint position controller is used during deployment.